# M. C. Rodriguez

### v1

@@ Abstract
Spontaneous breaking of a global symmetry in the Minimum Supersymmetric $SU(3)_{C}\otimes SU(3)_{L}\otimes U(1)_{N}$ Model.
@@ 1 Introduction
The Standard Model (SM) [1, 2], we introduce three families of leptons
@@
in parenthesis it appears the transformations properties under the respective factors $(SU(3)_{C},SU(2)_{L},U(1)_{Y})$, and $i=1,2,3,$ is family index. We want to emphasize, we only introduce left-handed neutrinos in agreement with the two component theory of neutrino theory of Landau [3], Lee and Yang [4] and Salam [5] from 1957, one very nice review about the history of SM can be found in [6].
In this model we have the following scalar field $H$
$H$ $=$ $\left(\begin{array}[]{c}h^{+}\\ h^{0}\end{array}\right)\sim({\bf 1},{\bf 2},+1).$ (6)
@@
$<H>$ $=$ $\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}0\\ v\end{array}\right),$ (9)
@@
wheree $\phi_{1}$ is CP even field and it get the following mass
$M^{2}_{H}=\sqrt{2}\mu,$ (13)
we must emphasize the SM do not predict any value for this mass, $\phi_{2}$ is CP odd field and it as massless field (Goldstone particles)  [^1] A nice review about it can be found in [7] and will be the responsable to give mass for $Z$ gauge boson, this mechanism is know as Brout-Englert-Higgs mechanism [7, 8]. After this break of symmetry we get massive gauge bosons
$M^{2}_{W}=\frac{g^{2}v^{2}}{4},\,\ M^{2}_{Z}=\frac{g^{2}v^{2}}{2\cos^{2}\theta_{W}},$ (14)
@@
The charged leptons get those masses from the following Yukawa coupling  [^2] For the masses of quarks, see [9, 10, 11].
@@
There are two main new results in particle physics, the first one is the Fermilab’s CDF collaboration for the mass of the $W$-boson mass [13]
@@
and the second one is the mass of the light Higgs boson at the LHC and its value is given by [14]
$M_{H}=125.09\pm 0.21\,\mathrm{(stat.)}\pm 0.11\,\mathrm{(syst.)}~{}\mathrm{GeV},$ (21)
the impact of those measurements in the new physics beyond SM can be found in [15]. Today we know neutrinos are massive particles [16, 17, 18, 19]. One possible way for generate mass for neutrinos is to break the residual lepton number conservation.
@@
$\Delta=\left(\begin{array}[]{cc}\Delta^{0}&\frac{h^{+}}{\sqrt{2}}\\ \frac{h^{+}}{\sqrt{2}}&H^{++}\end{array}\right)\sim\left({\bf 1},{\bf 3},2\right),\,\ \langle\Delta\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}v_{\rm T}&0\\ 0&0\end{array}\right)$ (22)
@@
$=$ $g^{\nu}_{ij}\left[-\overline{\nu^{c}_{iL}}\nu_{jL}\Delta^{0}+\left(\overline{l^{c}_{iL}}\nu_{jL}-\overline{\nu^{c}_{iL}}l_{jL}\right)\frac{h^{+}}{\sqrt{2}}+\overline{l^{c}_{iL}}l_{jL}H^{++}\right],$
@@
$M_{\rm M}\bar{\nu}^{c}_{L}\nu_{L},\,\ M_{\rm M}=g^{\nu}\frac{v_{\rm T}}{\sqrt{2}}e^{\imath\pi},$ (24)
it is know as Gelmini-Roncadelli scheme [20]. In order to obtain the positive mass, the eigenstates need to be redefined by the chiral rotations. The mass generation mechanism of charged leptons remains the same as presented in the SM, see our Eq.(16).
In this case the new scalar potential is given by
$V(H,\Delta)$ $=$ $V(\phi)+\mu^{2}_{1}Tr(\Delta^{\dagger}\Delta)+\lambda_{1}H^{\dagger}[Tr(\Delta^{\dagger}\Delta)]^{2}+\lambda_{2}H^{\dagger}Tr(\Delta^{\dagger}\Delta)$ (25)
$+$ $\lambda_{3}Tr(\Delta^{\dagger}\Delta\Delta^{\dagger}\Delta)+\lambda_{4}Tr(H^{\dagger}\Delta^{\dagger}\Delta H),$
where $V(H)$ is defined in our Eq.(10). Now if we make the shiftes defined in our Eq.(12), together with
$\Delta^{0}=\frac{1}{\sqrt{2}}\left(v_{\Delta}+R_{\Delta}+\imath I_{\Delta}\right),$ (26)
we obtain, from our potential defined in Eq.(25), the following constraints
$t_{H}$ $=$ $v\left(\mu^{2}+\lambda v^{2}+\frac{\lambda_{2}}{2}v^{2}_{\Delta}+\frac{\lambda_{4}}{2}v^{2}_{\Delta}\right),$
$t_{\Delta}$ $=$ $v_{\Delta}\left(\mu^{2}+\lambda_{1}v^{2}_{\Delta}+\frac{\lambda_{2}}{2}v^{2}_{\Delta}+\frac{\lambda_{4}}{2}v^{2}_{\Delta}+\lambda_{3}v^{2}_{\Delta}\right).$ (27)
and in order to obtain the minimum for our potential we need to impose
$t_{H}=t_{\Delta}=0.$ (28)
The mass matrix for neutral CP even $M^{2}_{R}$ is
$M^{2}_{R}=\left(\begin{array}[]{cc}2\lambda v^{2}+\frac{t_{H}}{v}&(\lambda_{2}+\lambda_{4})vv_{\Delta}\\ (\lambda_{2}+\lambda_{4})vv_{\Delta}&2(\lambda_{1}+\lambda_{3})v^{2}_{\Delta}+\frac{t_{\Delta}}{v_{\Delta}}\end{array}\right),$ (29)
the eigenvalues of this matrix are
$M^{2}_{H_{1}}$ $=$ $2\lambda v^{2}+2(\lambda_{1}+\lambda_{3})v^{2}_{\Delta}-v^{2}_{\Delta}\frac{4\lambda(\lambda_{1}+\lambda_{3})-(\lambda_{2}+\lambda_{4})^{2}}{2\lambda+2(\lambda_{1}+\lambda_{3})\frac{v^{2}_{\Delta}}{v^{2}}},$
$M^{2}_{H_{2}}$ $=$ $v^{2}_{\Delta}\frac{4\lambda(\lambda_{1}+\lambda_{3})-(\lambda_{2}+\lambda_{4})^{2}}{2\lambda+2(\lambda_{1}+\lambda_{3})\frac{v^{2}_{\Delta}}{v^{2}}}.$ (30)
While for CP odd $M^{2}_{I}$ we get
Therefore, as results we get We get two goldstone bosons, one will be eaten to give mass for $Z$-gauge boson and we get Majoron, $M$, which does not agree with the experimental width of $Z$-decay [21].
The masses of the gauge bosons in this case are given by [22, 23]
@@
therefore we can explain the CDF new results, given in our Eq.(20), if introduce a new real triplet Higgs boson [24, 25] and the vaccum expectation value, VEV, $v_{\rm T}$, of this new scalar must be of the order of $10$ GeV. Another interesting alternative is to introduce complex scalars and in this case the VEV must satisfy the following bound $v_{\rm T}\approx 3.2$ GeV [26].
@@
Supersymmetry, it is more known as SUSY, arose in theoretical papers more than 30 years ago independently by Golfand and Likhtman [27], Volkov and Akulov [28] and Wess and Zumino [29]. The supersymmetry algebra was introduced in [27]  [^3] They also constructed the first four-dimensional field theory with supersymmetry, (massive) quantum electrodynamics of spinors and scalars and as suggest we will call it as Golfand-Likhtman model and they show the original version in Russian of this article received by the Editorial Board on March, 10, 1971 [30]. The Volkov-Akulov article, however, deals only with fermions and its name was “Is the Neutrino a Goldstone Particle?”. For those interested in the history of Supersymmetry see [31, 32, 33, 34], while anyone interested, how to work with supersymmetry see the great review article [35].
@@
Another interesting possibility is to build Supersymmetric models with Higgs triplets as done in [40, 41, 42, 43]. There is also, an interesting proposal to explain this anomaly using the Minimal $R$-Symmetric extension of the Minimal Superymmetric Standard Model, known as MRSSM [44, 45, 46], and in this case we must respect the following inequality $|v_{\rm T}|\leq 4$ GeV [45].
@@
$SU(3)_{C}\otimes SU(3)_{L}\otimes U(1)_{N}$ (33)
are known as $3-3-1$ for short. They are interesting possibilities for the physics at the TeV scale [47, 48, 49].
@@
The minimal supersymmetric $SU(3)_{C}\otimes SU(3)_{L}\otimes U(1)_{N}$ model, or it is more known as MSUSY331 for shortness, [53, 54, 55, 56, 57]. We have have shown the lighest higgs has mass around 125.5 GeV [55, 57]. In an recente article, we have shown we can explain the $W$-boson mass presented by the CDF in the MSUSY331 [58]. All our previous analyses in the scalar potential, it was considered only the fields $\sigma^{0}_{2}$ and $\sigma^{\prime 0}_{2}$ get VEV. Now, we want to extend those analyses presented on [21] for the MSUSY331, it meand we want to consider that both $\sigma^{0}_{1}$ and $\sigma^{\prime 0}_{1}$ get VEV in similar conditions as presented in [58].
2 Minimal Supersymmetric 3-3-1 model.
The usual scalars fields in the minimal 3-3-1 model are defined as $\eta,\rho,\chi$ and $S$ [48] and we can decompose those fields in $SU(2)\otimes U(1)$ representations in the following way [21, 59, 60, 61]:
@@
As we saide, we have shown we can explain the $W$-boson mass presented by the CDF in the MSUSY331 in two possible ways [58]
@@
    $v_{\sigma_{1}}=v_{\sigma^{\prime}_{1}}=0$ GeV and $v_{\sigma_{2}}=10\,\ {\rm GeV}$ and $v_{\sigma^{\prime}_{2}}=11.19\,\ {\rm GeV}$.
@@
    Several solution as we shown in our Tab.(1). We will use the following point for our numerical analyses $v_{\sigma_{1}}=4.89\,\ {\rm GeV},v_{\sigma_{2}}=10\,\ {\rm GeV},v_{\sigma^{\prime}_{1}}=0.5\,\ {\rm GeV}$ and $v_{\sigma^{\prime}_{2}}=8.77\,\ {\rm GeV}$.

### v3

@@ Abstract
Spontaneous breaking of a global symmetry in the Minimum Supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ Model.
@@ 1 Introduction
In the context of the Standard Model (SM) [1, 2], we introduce only left-handed neutrinos in agreement with the two component theory of neutrino of Landau [3], Lee and Yang [4] and Salam [5] from 1957111About the history of SM can be found in [6] in the following represntations of $SU(2)_{L}$
@@
in parenthesis it appears the transformations properties under the respective factors $(SU(3)_{C},SU(2)_{L},U(1)_{Y})$, and $i=1,2,3,$ is family index. We are omitting the quark sector and they can be found at [7, 8, 9].
In this model to generate masses for the charged leptons, we have to introduce the following scalar field $H$ in the doublet representation
$\phi$ $=$ $\left(\begin{array}[]{c}h^{+}\\ h^{0}\end{array}\right)\sim({\bf 1},{\bf 2},+1).$ (6)
@@
$<\phi>$ $=$ $\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}0\\ v\end{array}\right),$ (9)
@@
under this expansion, our scalar potential become
$V(\phi)=V(h^{0})-2\mu^{2}\phi^{2}_{1}+hv\phi_{1}\left(\phi^{2}_{1}+\phi^{2}_{2}\right)+\frac{h}{4}\left(\phi^{2}_{1}+\phi^{2}_{2}\right)^{2},$ (13)
where $\phi_{1}$ is CP even field and it get the following mass [^2] A nice review about it can be found in [12].
$M^{2}_{H}=-2\mu^{2},$ (14)
we must emphasize the SM do not predict any value for this mass, in fact it can get any values rom 00 until any finite values, but we can get a superior limit in its mass given by $M_{H}<1000$ GeV [10, 11]. By another hand, the field $\phi_{2}$ is CP odd field and it as massless field (Goldstone particles) and will be the responsable to give mass for $Z$ gauge boson, this mechanism is know as Brout-Englert-Higgs mechanism [12, 13]. The mass of the light Higgs boson $h^{0}$ was measured at the LHC and its value is given by [14, 15]
$M_{H}=125.09\pm 0.21\,\mathrm{(stat.)}\pm 0.11\,\mathrm{(syst.)}~{}\mathrm{GeV}.$ (15)
After this break of symmetry we get massive gauge bosons throught the terms
$\left({\cal D}\phi\right)^{\dagger}\left({\cal D}\phi\right),$ (16)
where ${\cal D}$ is the covariant derivade of the higgs fields in the SM and we get the following results [7, 8]
$M^{2}_{W}=\frac{g^{2}v^{2}}{4}=\left(80.3505\pm 0.0077\right),\,\ M^{2}_{Z}=\frac{g^{2}v^{2}}{2\cos^{2}\theta_{W}},$ (17)
@@
The average values for the gauge bosnos masses are [14, 16]
$\left(M_{W}\right)_{ave}$ $=$ $\left(80.4133\pm 0.0080\right){\mbox{G}eV},$
$\left(M_{Z}\right)_{ave}$ $=$ $\left(91.1875\pm 0.0021\right){\mbox{G}eV},$ (19)
while for the Weinberg angle we have
$\sin^{2}\theta_{W}$ $=$ $0.2324\pm 0.0012$ (20)
There are also another interesting parameter to put some constrain in new physics, it is the $\rho$ parameter and it is define in the following way
$\rho\equiv\left(\frac{M^{2}_{W}}{M^{2}_{Z}\cos^{2}\theta_{W}}\right),$ (21)
in the SM its values is $1$ and it is in agreement with its experimental values given by
$\rho_{\rm EXP}=0.9998\pm 0.0008.$ (22)
The charged leptons get those masses from the following Yukawa coupling
@@
There is the Fermilab’s CDF collaboration for the mass of the $W$-boson mass [18]
@@
and the impact of this measurement together with Eq.(15) in the new physics beyond SM can be found in [19]. Today we know neutrinos are massive particles [20, 21, 22, 23]. One possible way for generate mass for neutrinos is to break the residual lepton number conservation.
@@
$\Delta=\left(\begin{array}[]{cc}\Delta^{0}&\frac{h^{+}}{\sqrt{2}}\\ \frac{h^{+}}{\sqrt{2}}&H^{++}\end{array}\right)\sim\left({\bf 1},{\bf 3},2\right),\,\ \langle\Delta\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}v_{\Delta}&0\\ 0&0\end{array}\right)$ (28)
@@
$=$ $g^{\nu}_{ij}\left[-\overline{\nu^{c}_{iL}}\nu_{jL}\Delta^{0}+\left(\overline{l^{c}_{iL}}\nu_{jL}-\overline{\nu^{c}_{iL}}l_{jL}\right)\frac{h^{+}}{\sqrt{2}}+\overline{l^{c}_{iL}}l_{jL}H^{++}+hc\right],$
@@
$M_{\rm M}\bar{\nu}^{c}_{L}\nu_{L},\,\ M_{\rm M}=g^{\nu}\frac{v_{\Delta}}{\sqrt{2}},$ (30)
it is know as Gelmini-Roncadelli scheme [24]. The mass generation mechanism of charged leptons remains the same as presented in the SM.
The masses of the gauge bosons in this case are given by [25, 26, 27]
@@
therefore we can explain the CDF new results, given in our Eq.(27), if introduce a new real triplet Higgs boson, as we shown in our Fig.(1), if $v_{\Delta}\approx 10$ GeV.
Figure 1: We show the predections for the gauge boson mass $W$ in the scheme of Gelmini-Roncadelli as a function of the Triplet VEV $v_{\Delta}$ and we show that we can explain the data from Eqs.(17,19,27).
By another hand the $\rho$ parameter in this case is given by [28]
$\rho$ $\simeq$ $1-2R\,\ R=\frac{v^{2}_{\Delta}}{v^{2}},$ (32)
and in this case the VEV must satisfy the following bound $v_{\Delta}\approx 4.92$ GeV, see our Fig.(2).
Figure 2: We show the predections for the parameter $\rho$, see our Eq.(22), in the scheme of Gelmini-Roncadelli as a function of the Triplet VEV $v_{\Delta}$.
In this scheme the scalar potential is given by [29, 30]
$V_{G}(\phi,\Delta)$ $=$ $\mu^{2}(\phi^{\dagger}\phi)+\mu^{2}_{1}Tr(\Delta^{\dagger}\Delta)+\lambda_{1}(\phi^{\dagger}\phi)^{2}+\lambda_{2}Tr(\Delta^{\dagger}\Delta)^{2}$
$+$ $\lambda_{3}(\phi^{\dagger}\phi)Tr(\Delta^{\dagger}\Delta)+\imath\lambda_{4}\epsilon_{ijk}(\phi^{\dagger}\sigma^{i}\phi)(\Delta^{j\dagger}\Delta^{k})+\lambda_{5}Tr(\Delta^{\dagger}\Delta\Delta^{\dagger}\Delta).$
Now if we make the shiftes defined in our Eq.(12), together with
$\Delta^{0}=\frac{1}{\sqrt{2}}\left(v_{\Delta}+R_{\Delta}+\imath I_{\Delta}\right).$ (34)
We get the following constraint equations
$t_{H}$ $=$ $v\left(\mu^{2}+\lambda_{1}v^{2}+\frac{\lambda_{3}}{2}v^{2}_{\Delta}-\frac{\lambda_{4}}{2}v^{2}_{\Delta}\right)=0,$
$t_{\Delta}$ $=$ $v_{\Delta}\left(\mu^{2}_{1}+\lambda_{2}v^{2}_{\Delta}+\frac{\lambda_{3}}{2}v^{2}_{\Delta}-\frac{\lambda_{4}}{2}v^{2}_{\Delta}+\lambda_{5}v^{2}_{\Delta}\right)=0.$ (35)
The mass spectrum of the neutral scalar is
$V_{quad}$ $=$ $\frac{1}{2}\left(\begin{array}[]{cc}R_{H}&R_{\Delta}\end{array}\right)M^{2}_{R}\left(\begin{array}[]{c}R_{H}\\ R_{\Delta}\end{array}\right)+\frac{1}{2}\left(\begin{array}[]{cc}I_{H}&I_{\Delta}\end{array}\right)M^{2}_{I}\left(\begin{array}[]{c}I_{H}\\ I_{\Delta}\end{array}\right).$ (42)
For the scalar of CP even, we get the following mass matrix
$M^{2}_{R}=\left(\begin{array}[]{cc}2\lambda_{1}v^{2}+\frac{t_{H}}{v}&(\lambda_{3}-\lambda_{4})vv_{\Delta}\\ (\lambda_{3}-\lambda_{4})vv_{\Delta}&2\lambda_{2}v^{2}_{\Delta}+\frac{t_{\Delta}}{v_{\Delta}}\end{array}\right).$ (43)
We get two massives scalar field, the light one we will call as $h^{0}$ and the heavier as $H^{0}$. The mass of the light scalar $h$ is [^3] The mass of heavy higgs $H$ we will present below.
$m^{2}_{h}\approx 2\lambda_{1}v^{2},$ (44)
the mass $m_{h}$ as function of the VEV $v_{\Sigma}$ is shown in our Fig.(3), where we can explain the experimental. For the case of CP odd sector the mass matrix has the following representation
We get two goldstone bosons, one is eaten by $Z^{0}$ and the second is call as Majoron $M$ is $I_{\Delta}$ [24] and it is inconsistent with the experimental data [31].
Figure 3: We show the predections for the light higgs mass $h$ in the scheme of Gelmini-Roncadelli as a function of the parameter $\lambda_{1}$ and we show that we can explain the data from Eq.(15).
In this scheme the scalar potential is given by [29, 30]
$V(\phi,\Delta)$ $=$ $V_{G}(\phi,\Delta)+\mu_{M}\phi^{\dagger}\Delta\tilde{\phi},$ (46)
where $V_{G}(\phi,\Delta)$ is defined in our Eq.(LABEL:potmajorongelmini) and $\tilde{\phi}\equiv\left(\imath\sigma_{2}\phi\right)$. Under this hipotesis our Eq.(47) become
$M^{2}_{R}=\left(\begin{array}[]{cc}2\lambda_{1}v^{2}&(\lambda_{3}-\lambda_{4})vv_{\Delta}+2\sqrt{2}\mu_{M}v\\ (\lambda_{3}-\lambda_{4})vv_{\Delta}+2\sqrt{2}\mu_{M}v&2\lambda_{2}v^{2}_{\Delta}-\sqrt{2}\mu_{M}\frac{v^{2}}{v_{\Delta}}\end{array}\right).$ (47)
and the mass spectrum, beyond $h$ given in Eq.(44), we get the heavy huggs $H$ and its mass is
$m^{2}_{H}\approx-\sqrt{2}\mu_{M}\frac{v^{2}}{v_{\Delta}},$ (48)
$m_{H}$ as function of the VEV $v_{\Delta}$ is shown in our Fig.(4).
Figure 4: We show the predections for the heavy Higgs mass $H$ in the scheme of Gelmini-Roncadelli, see Eq.(48), as a function of the Triplet Field VEV $V_{\Delta}$.
The mass matrix for neutral CP odd $M^{2}_{I}$ we get [29, 30]
$M^{2}_{I}=\left(\begin{array}[]{cc}-4\mu_{M}v_{\Delta}&2\mu_{M}v\\ 2\mu_{M}v&-\mu_{M}\frac{v}{v_{\Delta}}\end{array}\right).$ (49)
the majron mass
$m^{2}_{M}\approx-\sqrt{2}\mu_{M}\left(\frac{4v^{2}_{\Delta}+v^{2}}{v_{\Delta}}\right),$ (50)
$m_{M}$ as function of the VEV $v_{\Delta}$ is shown in our Fig.(5).
Figure 5: We show the predections for the Majoron mass $M$, see Eq.(50), in the scheme of Ma as a function of the Triplet Field VEV $V_{\Delta}$.
@@
Supersymmetry, it is more known as SUSY, arose in theoretical papers more than 30 years ago independently by Golfand and Likhtman [32], Volkov and Akulov [33] and Wess and Zumino [34]. The supersymmetry algebra was introduced in [32]  [^4] They also constructed the first four-dimensional field theory with supersymmetry, (massive) quantum electrodynamics (SQED) of spinors and scalars and as suggest we will call it as Golfand-Likhtman model and they show the original version in Russian of this article received by the Editorial Board on March, 10, 1971 [35].. The Volkov-Akulov article, however, deals only with fermions and its name was “Is the Neutrino a Goldstone Particle?”. For those interested in the history of Supersymmetry see [36, 37] and for the earliest supersymmetric Standard Model see [38, 39], while anyone interested, how to work with supersymmetry see the great review article [40].
@@
Another interesting possibility is to build Supersymmetric models with Higgs triplets as done in [45, 46, 47, 48]. There is also, an interesting proposal to explain this anomaly using the Minimal $R$-Symmetric extension of the Minimal Superymmetric Standard Model, known as MRSSM [49, 50, 51], and in this case we must respect the following inequality $|v_{\Delta}|\leq 4$ GeV [50].
@@
$SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ (51)
are known as $331$ for short. They are interesting possibilities for the physics at the TeV scale [52, 53, 54].
@@
The minimal supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ model, or it is more known as MSUSY331 for shortness, [58, 59, 60, 61, 62]. We have have shown the lighest higgs has mass around 125.5 GeV [60, 62]. In an recente article, we have shown we can explain the $W$-boson mass presented by the CDF in the MSUSY331 [28]. All our previous analyses in the scalar potential, it was considered only the fields $\sigma^{0}_{2}$ and $\sigma^{\prime 0}_{2}$ get VEV. Now, we want to extend those analyses presented on [25] for the MSUSY331, it meand we want to consider that both $\sigma^{0}_{1}$ and $\sigma^{\prime 0}_{1}$ get VEV in similar conditions as presented in [28].
2 Minimal Supersymmetric $331$ model.
The usual scalars fields in the minimal $331$ model are defined as $\eta,\rho,\chi$ and $S$ [53] and we can decompose those fields in $SU(2)\times U(1)$ representations in the following way [25, 63, 64, 65]:
@@
As we said, we have shown we can explain the $W$-boson mass presented by the CDF in the MSUSY331 in two possible ways [28]. For somephenomenological analyses is useful define the following parameter [68, 69] [^6] We want to perform similar constraints as presented in [68, 69] in this model.
$\xi_{WW^{\prime}}=\left(\frac{M_{W}}{M_{W^{\prime}}}\right)^{2},\,\ \xi_{ZZ^{\prime}}=\left(\frac{M_{Z}}{M_{Z^{\prime}}}\right)^{2}$ (99)
@@
    $v_{\sigma_{1}}=v_{\sigma^{\prime}_{1}}=0$ GeV and $v_{\sigma_{2}}=10\,\ {\rm GeV}$ and $v_{\sigma^{\prime}_{2}}=11.19\,\ {\rm GeV}$ and the parameters defined in our Eq.(99) is shown in our Tab.(1).
    $v_{\chi}$ (GeV) $\xi_{WW^{\prime}}$ $\xi_{ZZ^{\prime}}$
    1000 $4.09\times 10^{-4}$ $2.03\times 10^{-3}$
    1500 $2.04\times 10^{-4}$ $1.25\times 10^{-3}$
    2000 $1.02\times 10^{-4}$ $8.14\times 10^{-4}$
    2500 $5.42\times 10^{-5}$ $5.61\times 10^{-4}$
    3000 $3.04\times 10^{-5}$ $4.07\times 10^{-4}$
    3500 $1.81\times 10^{-5}$ $3.07\times 10^{-4}$
    4000 $1.13\times 10^{-5}$ $2.39\times 10^{-4}$
    4500 $7.45\times 10^{-6}$ $1.91\times 10^{-4}$
    5000 $5.06\times 10^{-6}$ $1.56\times 10^{-4}$
    Table 1: The parameters $\xi_{WW^{\prime}}$ and $\xi_{ZZ^{\prime}}$, see Eq.(99), as function of $v_{\chi}$ in GeV and we take $v_{\chi^{\prime}}=1000$ GeV and the others vev are given in Eqs.(63,64) and Eq.(73) [28].
@@
    Several solution as we shown in our Tab.(2). We will use the following point for our numerical analyses $v_{\sigma_{1}}=4.89\,\ {\rm GeV},v_{\sigma_{2}}=10\,\ {\rm GeV},v_{\sigma^{\prime}_{1}}=0.5\,\ {\rm GeV}$ and $v_{\sigma^{\prime}_{2}}=8.77\,\ {\rm GeV}$ and the parameters defined in our Eq.(99) is shown in our Tab.(3).
@@
$v_{\chi}$ (GeV) $\xi_{WW^{\prime}}$ $\xi_{ZZ^{\prime}}$
1000 $4.09\times 10^{-4}$ $2.03\times 10^{-3}$
1500 $2.04\times 10^{-4}$ $1.25\times 10^{-3}$
2000 $1.02\times 10^{-4}$ $8.14\times 10^{-4}$
2500 $5.42\times 10^{-5}$ $5.61\times 10^{-4}$
3000 $3.04\times 10^{-5}$ $4.07\times 10^{-4}$
3500 $1.81\times 10^{-5}$ $3.07\times 10^{-4}$
4000 $1.13\times 10^{-5}$ $2.39\times 10^{-4}$
4500 $7.45\times 10^{-6}$ $1.91\times 10^{-4}$
5000 $5.06\times 10^{-6}$ $1.56\times 10^{-4}$
Table 3: The parameters $\xi_{WW^{\prime}}$ and $\xi_{ZZ^{\prime}}$, see Eq.(99), as function of $v_{\chi}$ in GeV and we take $v_{\chi^{\prime}}=1000$ GeV and the others vev are given in Eqs.(63,64) and Eq.(73) [28].